APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Sunday September 3, 2006 (MJD 53982) ---=== OBSERVING TEAM ===--- Swing: Kaike Pan Night: Mike Harvanek Support: Steph Snedden: setup for the coming ET run ---=== OBSERVING PLAN ===--- SN Science! ---=== OBSERVING SUMMARY ===--- Clouds, rain and high humidity all night. Never opened. The alt fiducials at 75 and 90 degrees did not read in either direction. The Lindy cart is still in the enclosure due to rain. ---=== OBSERVING LOG ===--- Afternoon: ---------- The PT is back to operational status. Night: ------ The alt fiducials at 75 and 90 degrees did not read in either direction during telescope checkout. Clouds, rain and high humidity kept us closed all night. The Lindy cart is still in the enclosure due to the rain. endNight rejoiced. ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- 6301 23:26Z 23:53Z 100 O 43.72 50.32 46 ignore ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): N/A QA Procedures Done (y/n): N/A UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== DATA TAPE SUMMARY ===--- Goes: JL Stays: JL ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 23:21Z 47 48 14 199 (SSW) 76073 - clouds/rain 23:52Z 48 49 13 184 (S) 31684 - clouds/rain 00:22Z 48 49 9 187 (S) 25679 - clouds/rain 00:54Z 49 49 4 183 (S) 40016 - clouds/rain 02:01Z 47 47 4 81 (E) 88 - clouds/rain 02:33Z 46 47 9 124 (SE) 96 - clouds/rain 03:05Z 46 47 8 107 (ESE) 74 - clouds/rain 03:37Z 46 47 9 122 (ESE) 116 - clouds/rain 04:10Z 47 47 10 122 (ESE) 118 - clouds/rain 04:42Z 47 47 13 127 (SE) 106 - clouds/rain 05:14Z 47 47 7 145 (SE) 127 - clouds/rain 05:46Z 47 47 7 134 (SE) 207 - clouds/rain 06:18Z 48 48 13 146 (SE) 26336 - clouds/rain 06:50Z 48 48 9 128 (SE) 14895 - clouds/rain 07:23Z 48 48 10 143 (SE) 12409 - clouds/rain 07:55Z 48 48 8 173 (S) 51581 - clouds/rain 08:27Z 48 48 8 177 (S) 57965 - clouds/rain 08:59Z 48 48 12 200 (SSW) 82059 - clouds/rain 09:31Z 48 48 6 185 (S) 40205 - clouds/rain 10:03Z 48 48 8 145 (SE) 33718 - clouds/rain ---=== TELESCOPE STATUS ===--- Never opened due to clouds. rain and high humidity. Telescope stowed at 30 degrees with cartridge 1 (plate 8021) mounted. Counterweights at 250. Spectro auto-fill connected. Spectro dewar 10:02Z 132 lb, 25 psi Imager dewar 10:02Z 262 lb, 25 psi No interlocks bypassed. ---=== SOFTWARE USED ===--- IOP/SOP: v4_26_0 Watcher: v2_34_0 MCP: v6_3_2 TPM: tpm_v3_3_0 AstroDa: v15_12_1 TCC: TCC 2.7.2.1 August 6 2004 sdssProcedures: v2_19 SoS: v5_1_4 hoggPT: v1_6_9 plate-mapper: v4_3_1 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A 24.6740 24.6750 Axial B 26.2530 26.2540 Axial C 25.3480 25.3490 Trans D 13.5740 13.5760 Lateral E 9.9568 9.9568 Lateral F 11.5316 11.5316 GALILS Commanded: 4600. -2000. 2650. -9300. -9800. -9800. Actual: 4600. -2000. 2650. -9300. -9800. -9800. SETMIR VALUES PriDesOrient: 0.00 0.00 0.00 0.00 0.00 PriOrient: 0.00 0.00 0.00 0.00 0.00 SECONDARY: ---------- Focus: 0.00 Air Temp.: 9.2 Alt.: 30.036167 MIGS TONIGHT NOMINAL Axial A 1.5570 1.5550 Axial B 1.1330 1.1290 Axial C 1.1940 1.1910 Trans D -0.0600 -0.0530 GALILS Commanded: 1614068. 1594704. 1543101. -10550. -5800. Actual: 1614050. 1594700. 1543100. -10550. -5800. SETMIR VALUES SecDesOrient: 1257.00 20.00 -10.00 -100.00 176.64 SecOrient: 1257.00 19.99 -10.01 -100.44 176.56 ---=== PROBLEMS IN DETAIL ===---